OpenStudy (anonymous):

graph the equation and identify the y intercept x+2y=6

OpenStudy (saifoo.khan):

Mini wanna try?

OpenStudy (mimi_x3):

y-intercept: x= 0 0 + 2y = 6 2y = 6 y=3 Am i right ? it has been a while that i have done these type of questions xD

OpenStudy (saifoo.khan):

x intercept is 6. >_<

OpenStudy (saifoo.khan):

x = (6,0) y = (0,3)

OpenStudy (mimi_x3):

How ?

OpenStudy (saifoo.khan):

When y intercepts x = 0 when x intercepts y = 0

OpenStudy (saifoo.khan):

x+2y=6 x + 2(0) = 6 x = 6 (6,0)

OpenStudy (mimi_x3):

I am right, i was finding the y-intercept. The question says y-intercept if you're not blind :P

OpenStudy (saifoo.khan):

u have to graph the equation, for that u need both things. :P

OpenStudy (mimi_x3):

I know, but you were like im "incorrect" :P

OpenStudy (saifoo.khan):

where?

OpenStudy (mimi_x3):



OpenStudy (saifoo.khan):